# Explainable Artificial Intelligence techniques for interpretation of food datasets: a review

### v1

@@ 10 Open Challenges and Future Directions
Through a comparative analysis of various studies and an evaluation of their statistical insights, we have identified several opportunities for further research and future XAI applications in the food quality field. Encouraging their application to existing studies is essential for validating findings and ensuring transparency. Notably, the expanding range of XAI methods presents promising opportunities for analyzing underexplored data types. Spectral data, commonly used in physicochemical analysis, has yet to be adequately addressed by current XAI techniques. Closing this gap will require the development of tailored approaches to interpret this complex data type. Additionally, there are relatively few XAI methods specifically designed to explain AI models used in time series analysis, as confirmed by [218]. Researchers should stay attentive to new developments, as applying and experimenting with emerging techniques helps to refine and promote their use.
@@
According to [103] and to the works summarized by [5], the explanations provided by many XAI methods are often difficult to interpret and require the expertise of a field specialist, turning the explanations themself into additional steps to be deciphered. To address this, recent advancements in XAI research propose new solutions, such as the development of frameworks, metrics to evaluate model outputs [219, 220], and the use of generative AI to simplify and clarify the explanations provided by existing methods. It is crucial to utilize multiple methods from those available to gain a comprehensive understanding of a model’s decision-making process. Specifically, by adopting a glocal approach [5], which combines both local and global explanations of the same model, more complete and user-friendly explanations can be achieved. By studying the work of [218], we can observe that, to our knowledge, several types of XAI techniques have not been utilized in the analyzed works, yet they could prove extremely useful in various contexts. For instance, rule-based techniques, which we appreciated in just some studies [31], offer the potential to uncover causal relationships between the physicochemical properties of food products, adding depth to the interpretation of these behaviours and aiding in the development of more interpretable models. Concept-based learning algorithms represent a popular category of methods that can be used to explain model predictions in terms of adjectives, concepts, or abstractions easily understood by humans [221].
Another growing approach in food quality assessments is data fusion, as it integrates multiple types of data, such as chemical, physical, and sensory information, to make more comprehensive decisions about food products. This fusion of diverse data types enables a richer analysis but makes understanding the outcomes more challenging. To address this complexity, hierarchical-based explainability approaches could be proposed to break down the contribution of each data type to the final decision. Although not yet well-defined, this type of XAI could offer a viable solution for explaining models that integrate multiple data types. By introducing a hierarchy within explanations, it becomes possible to discern the contribution of each data type to the overall result, clarifying how the combined dataset influences the model’s decisions.
@@
Counterfactual explanations, extensively studied in XAI research [5], can significantly support in silico food simulation research. These techniques help to understand how variations in ingredients, environmental conditions, and processes impact food quality, taste, or nutritional profile. Counterfactual explanations also allow researchers to examine how specific ingredient changes might influence shelf life. By simulating alternative pathways without requiring costly or time-intensive experiments, counterfactual methods can guide decision-making, optimize formulations, and enhance the accuracy of outcome predictions.
@@ 11 Conclusion
EXplainable Artificial Intelligence (XAI) techniques have emerged as important tools for enhancing the transparency, trustworthiness, and auditability of AI models, supporting the production of reliable and understandable outcomes. These requirements are essential in food engineering, as food is a fundamental aspect of human life and its quality and safety need to be studied with careful attention. In this survey, we aimed to bridge the gap between these two disciplines by emphasizing the importance of XAI techniques and offering practical insights into both domains. Our comprehensive review examined a wide array of studies from the literature, which we categorized according to the types of data utilized—tabular, pictorial, spectral, and time series—and the forms of explainability provided, including numerical, rule-based, visual, and mixed explanations. Additionally, we proposed a food quality taxonomy to contextualize the research, focusing on key areas such as food safety, nutritional value, sensory attributes, authenticity and traceability, and sustainability and healthiness. Finally, we conducted a comparison of the studies to uncover valuable insights, identifying main trends, strengths, and divergences in the current research landscape. This analysis allowed us to pinpoint critical areas where XAI can drive advancements in food quality.
@@ Acknowledgments
12 Declaration of generative AI and AI-assisted technologies in the writing process

### v2

@@ 10 Open Challenges and Future Directions
Based on the analysis presented in this survey, two key directions emerge for improving the food industry’s supply chain through the integration of AI. The first direction arises from the growing prevalence of AI technologies in the world and the resulting adaptation by both the industry and society. This is a trend that the food industry should also embrace. The second direction focuses on the development of new technologies in the field of XAI. These technologies have the potential to enhance the application of AI within the food industry and expand its possible uses. In the following subsections, we will discuss these two directions in more detail.
To provide a concise roadmap, we summarize the five principal research gaps emerging from the surveyed literature.
  • (G1)
    Real-time interpretable imaging at scale.
  • (G2)
    Actionable counterfactual explanations for process and formulation decisions.
  • (G3)
    Human-in-the-loop deployment and validation protocols.
  • (G4)
    Benchmarks and formal evaluation metrics for explanation quality.
  • (G5)
    Regulatory compliance, traceability, and post-deployment monitoring for high-stakes applications.
These gaps are revisited throughout the following subsections and collectively define a prioritized set of directions for both future research and industrial adoption.
10.1 Integrating AI into the Food Engineering
Despite the increasing interest in interpretable modeling, the adoption of XAI within food engineering remains relatively low. As discussed by [137] and evidenced throughout the surveyed literature, this limited uptake can be attributed to several interrelated factors.
Some limitations are primarily practical. First, the effective application of XAI methods requires multidisciplinary expertise, combining data science proficiency or deep domain knowledge in food technology, which is still scarce across research groups and industrial teams. Moreover, the computational and implementation costs associated with advanced interpretability frameworks—such as SHAP, Grad-CAM, or hybrid visualization-based pipelines—remain prohibitive for many laboratories and small-scale producers, particularly when dealing with high-dimensional spectral or multimodal datasets. Additionally, there is an overall lack of formal evaluation of explanations, a topic that poses several challenges, mainly due to the missing—or difficult to obtain—ground truths. However, as noted by [160], it is paramount that the quality of explanations be evaluated in a formal way, and not only in an anecdotal manner or via user studies.
From a deployment standpoint, a critical bottleneck is represented by (G1) real-time interpretable imaging at scale. Many food industry scenarios rely on high-throughput vision systems for tasks such as defect detection, contamination screening, and product grading, where explanations must be generated under strict latency constraints and remain stable across varying acquisition conditions, including illumination, camera geometry, and packaging reflections, as well as across seasonal and batch-to-batch changes. Achieving low-latency and robust explanations for imaging models remains an open engineering and research challenge, particularly when transitioning from controlled benchmark datasets to continuously evolving production environments.
The regulatory landscape of the food sector also imposes barriers to the use of AI. Strict safety standards and validation requirements often discourage the deployment of novel AI-driven decision systems unless their traceability and reliability can be fully demonstrated [137, 51, 56]. On the side of AI-specific regulations, the EU AI Act [2] mandates transparency as a fundamental property of High-risk AI systems. Specifically, in the food sector, [236] argues that many AI applications could fall under the High-risk umbrella due to their direct impact on human health or other societal impacts. While it could be argued that transparency is not necessarily solely connected with explainability [172], XAI could still be used as a tool to enhance human oversight of these systems. The extent to which XAI provides a quantifiable contribution to transparency, as mandated by the EU AI Act, is to be determined by conformity assessments. The regulation mandates that these be carried out, possibly by third-party entities, to determine whether these systems comply with the Act. Thus, any consideration of guidelines for XAI in food engineering will have to be determined on a case-by-case basis.
This directly connects to (G5) regulatory compliance, traceability, and auditing. When artificial intelligence systems influence decisions pertaining to food safety, product labeling, or fraud detection, stakeholders require traceable evidence linking model outputs to documented decision criteria, in a form that can support inspection, reporting, and conformity assessments. As a consequence, XAI should not be treated merely as a post-hoc visualization layer, but rather as an integral component of a broader compliance workflow that encompasses documentation, model versioning, and repeatable explanation generation under controlled and standardized conditions.
Remaining in the context of AI regulations, the US Food and Drug Administration (FDA) also mandates requirements for the usage of AI systems, albeit for the development of “drugs and biological products[235]. Similar to the EU AI Act, these considerations propose a framework based on risk and context of use to determine the level of transparency of the system. This could translate seamlessly to food applications, given, as previously stated, the potential for high risk in some of the applications surveyed in the present work. In addition to this risk-based framework, the FDA also emphasizes the need to monitor the AI system after deployment. XAI techniques, in this context, have been shown to help identify post-deployment phenomena such as distribution shifts [10] and concept drift [115, 228].
Post-deployment monitoring is also tightly coupled with (G3) human-in-the-loop deployment and validation. In real production environments, explanations must be reviewed and acted upon by inspectors and quality managers, and the system should support structured feedback loops encompassing the flagging of systematic explanation drift, the triggering of re-calibration procedures, and the activation of escalation protocols when anomalous behavior is detected. Designing reliable operational workflows that integrate automated alarms with human verification, therefore, remains a central requirement for the reliable adoption of XAI in industrial food quality control.
In conclusion, while AI has become an increasingly essential tool throughout the food supply chain, its use remains intrinsically limited by its opaque nature. XAI, which offers an effective response to the black-box problem by providing clarity and transparency, also presents significant challenges. Due to this complexity, research in the field of XAI is highly active and rapidly evolving. Existing techniques are continually being updated and improved, while new methods are consistently emerging. Researchers should stay informed about these advancements, as applying and experimentally evaluating new techniques can help refine them and encourage their adoption.
10.2 Advancing XAI Technologies for the Food Supply Chain
Through the comparative analysis of the studies surveyed and an evaluation of their statistical insights, we have identified several opportunities for further research and future XAI applications in the field of food quality.
By studying the work of [27], we can observe that, to our knowledge, several types of XAI techniques have not been utilized in the analyzed works, yet they could prove extremely useful in various contexts. For instance, rule-based techniques, which we appreciated in just some studies [193, 196, 234, 77, 212], offer the potential to uncover causal relationships between the physicochemical properties of food products, adding depth to the interpretation of these behaviours and aiding in the development of more interpretable models.
@@ 10.2 Advancing XAI Technologies for the Food Supply Chain
We can also observe that, according to [45] and to the works summarized by [124], the explanations provided by many XAI methods are often difficult to interpret and require the expertise of a field specialist, turning the explanations themself into additional steps to be deciphered. To address this, recent advancements in XAI research propose new solutions, such as the development of frameworks, metrics to evaluate model outputs [11, 160], and the use of generative AI to simplify and clarify the explanations provided by existing methods, e.g., by using Large Language Models to provide text-based explanations that helps interpreting explanations [218]. Moreover, utilizing various methods helps achieve a thorough understanding of a model’s decision-making process. Specifically, by adopting a glocal approach [124], which combines both local and global explanations of the same model, more complete and user-friendly explanations can be achieved.
Among the various new techniques for which no applications were identified in this survey, concept-based learning, algorithms represent a popular category of methods that can be used to explain model predictions in terms of adjectives, concepts, or abstractions easily understood by humans [107].
@@
Additionally, counterfactual explanations, extensively studied in XAI research [124], can significantly support in silico food simulation research. These techniques help to understand how variations in ingredients, environmental conditions, and processes impact food quality, taste, or nutritional profile. Counterfactual explanations also allow researchers to examine how specific ingredient changes might influence shelf life. By simulating alternative pathways without requiring costly or time-intensive experiments, counterfactual methods can guide decision-making, optimize formulations, and enhance the accuracy of outcome predictions.
This motivates the need for (G2) actionable counterfactual explanations. Beyond answering why a particular decision was reached, many food engineering use cases require guidance on what should change in order to achieve a desired outcome, such as improving a quality grade, extending shelf life, or reducing defect probability. Future work should therefore focus on counterfactual generation that is (i) feasible under process constraints, including bounded temperature ranges, permitted additives, and production tolerances, (ii) consistent with the interdependencies among correlated variables such as moisture and texture interactions, and (iii) validated against domain knowledge and, where possible, confirmatory experimental evidence.
Another potential area for improvement is the use of hybrid systems that combine black-box models with white-box components, as seen in Neurosymbolic AI [25]. Another strategy could involve training black-box models on engineered features rather than on raw data, a common approach in time-series analysis [108]. This method may enhance interpretability for experts in the field.
An alternative emerging approach in food quality assessments is data fusion, as it integrates multiple types of data, such as chemical, physical, and sensory information, to make more comprehensive decisions about food products. This fusion of diverse data types enables a richer analysis but makes understanding the outcomes more challenging. To address this complexity, hierarchical-based explainability approaches could be proposed to break down the contribution of each data type to the final decision. Although not yet well-defined, this type of XAI could offer a viable solution for explaining models that integrate multiple data types. By introducing a hierarchy within explanations, it becomes possible to discern the contribution of each data type to the overall result, clarifying how the combined dataset influences the model’s decisions.
In summary, these approaches provide ways to enhance the reliability and transparency of AI algorithms across different areas of the food industry. As highlighted by [151], XAI should be implemented in human-in-the-loop systems primarily as diagnostic tools. This integration can help improve human reliability by allowing users to identify mistakes or biases in the relationships learned by these systems. However, these tools offer limited support for real-time applications aimed at increasing efficiency. One potential application for XAI in real-time systems would be to randomly sample predictions from the model, generate explanations for those predictions, and have experts evaluate them. This process could help uncover prediction errors and prompt necessary interventions within the system.
This direction is closely aligned with (G3) human-in-the-loop deployment and (G4) benchmarks and formal evaluation metrics. Robust adoption requires both operational protocols and measurable quality criteria for assessing the validity of explanations. Beyond anecdotal examples, the research community needs shared benchmarks, standardized reporting of explainer configurations including baseline choices and perturbation strategies, and evaluation suites capturing key properties such as faithfulness, stability, and human usability. In food quality applications specifically, benchmark design should explicitly incorporate batch effects, seasonal variability, and acquisition shifts arising from sensor and imaging condition changes, to reflect realistic industrial distributions rather than static and controlled laboratory settings.
Finally, the expanding range of XAI methods offers promising opportunities to analyze underexplored data types. Spectral data, commonly used in physicochemical analysis, has yet to be adequately addressed by current XAI techniques. Closing this gap will require the development of tailored approaches to interpret this complex data type. Additionally, there are relatively few XAI methods specifically designed to explain AI models used in time series analysis, as confirmed by [27].
Overall, addressing (G1) through (G5) jointly is essential for advancing from explanatory prototypes toward reliable and deployable systems. Real-time explanation generation must be paired with actionable counterfactual guidance, human-centered validation loops, benchmark-driven evaluation, and compliance-ready documentation and monitoring frameworks. Progress on any single gap in isolation is unlikely to be sufficient; it is their coordinated resolution that will enable the reliable integration of XAI into operational food quality control pipelines.
@@ 11 Conclusion
EXplainable Artificial Intelligence (XAI) techniques have emerged as important tools for enhancing the transparency and auditability of AI models, supporting the production of reliable and understandable outcomes. These requirements are essential in food engineering, as food is a fundamental aspect of human life, and its quality and safety need to be studied with careful attention. In this survey, we aimed to bridge the gap between these two disciplines by emphasizing the importance of XAI techniques and offering practical insights into both domains.
This work targets researchers and professionals in food science and technology interested in using XAI to enhance model transparency and regulatory compliance in food applications. It also serves as a reference for AI developers and policymakers focused on creating explainable systems in quality control.
Our comprehensive review is, to our knowledge, the first to connect the fields of XAI and food by examining how explainability techniques can be applied to food quality. We analyzed a more than 200 studies from the literature and categorized them according to the types of data utilized—tabular, pictorial, spectral, and time series—and the forms of explainability provided, including numerical, rule-based, visual, and mixed explanations. Additionally, we proposed a food quality taxonomy to contextualize the research, focusing on key areas such as food safety, nutritional value, sensory attributes, authenticity and traceability, and sustainability and healthiness. This taxonomy aims to unify existing literature and provide a consistent overview of the current state of XAI applications in food quality analysis.
Finally, we conducted a comparison of the studies to uncover valuable insights, identifying main trends, strengths, and divergences in the current research landscape. This analysis enabled us to map the current use of XAI in food quality applications, highlighting both its well-established applications and the areas that require further exploration. In particular, two key directions emerge: the broader adoption of AI (and XAI) across the food supply chain, and the development of more advanced XAI methods to enhance the transparency and impact of these applications.
Statements and Declarations
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of generative AI and AI-assisted technologies in the writing process